# ./references.bib

### v1

@@ 3 Clifford composite operator
If the Clifford composite operator is singular then we are done and $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. So we will assume that the Clifford composite operator is invertible. We will use linearity to reduce to the case where $(\bm{\lambda},\nu)=\bm{0}$. If $\sqrt{\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}\right\rVert+\left\lVert B\bm{\psi}\right\rVert+\left\lVert B^{\dagger}\bm{\psi}\right\rVert}\leq\epsilon$ for some unit $\bm{\psi}$ we will define
$\bm{\Psi}=\frac{1}{\sqrt{2m}}\begin{bmatrix}\bm{\psi}\\ \vdots\\ \bm{\psi}\end{bmatrix}$
@@
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert\leq\frac{1}{\sqrt{2m}}\sqrt{\left(\sum_{i=1}^{d}2m\left\lVert A_{i}^{2}\bm{\psi}\right\rVert^{2}\right)+2m\left\lVert B^{\dagger}B\bm{\psi}\right\rVert^{2}+2m\left\lVert BB^{\dagger}\bm{\psi}\right\rVert^{2}}\leq\epsilon$
implies
$\leq\epsilon+\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert.$
Lastly note that since the Clifford composite operator is invertible,
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{2}=\left\lVert(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right))^{-1}\right\rVert.$
This implies that
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{-1}\right\rVert^{-1}$ $=\sqrt{\min_{\left\lVert\mathbf{v}\right\rVert=1}\left\lVert\left(L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\right)\mathbf{v}\right\rVert}$
$\leq\sqrt{\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert}.$
@@
The lack of a proof does not hinder us from proceeding with the generalization of the theory in [cerjan_loring_vides_2023]. The fact that for a normal matrix $A$,
@@
In any case we are working with a non-normal matrix, and with luck the matrix can be nearly Hermitian or nearly normal. However we do not work with luck and thus we need a way to characterize what we mean by nearly normal. We will make use of definition and terminology presented in [ruhe] in order to characterize deviation from normality. In particular we have,
@@
This is in some sense equivalent to considering the quantity $||AA^{\dagger}-A^{\dagger}A||$ and formal statements comparing the two quantities can be found in Theorem 2 of [elsner_paardekooper]. Also note that when the Clifford composite operator does not deviate too much from normal then we get some decent spectral variation properties.
@@ 4 Quadratic composite operator
In [cerjan_loring_vides_2023] and [loring], it was shown that when it comes to finding unit states that are approximate joint eigenstates to the obersables/operators, both the spectral localizer’s $\epsilon$-pseudospectrum and the quadratic’s $\epsilon$-pseudospectrum provide good methods in the all Hermitian observables setting. In Section 3 we have a non-Hermitian matrix/Hamiltonian in the construction of the Clifford composite operator in order to generalized the spectral localizer $\epsilon$-pseudospectrum theory. The appeal of the spectral localizer $\epsilon$-psedudospectrum in the Hermitian case and Clifford composite operator $\epsilon$-pseudospectrum in the non-Hermitian case is that is also provides information about a system’s topology see [cerjan_loring_pho_2022],[cerjan_loring_2022],[kahlil_cerjan_loring] for the Hermitian setting and [cerjan_koekebier_Schulz_Baldes] for the non-Hermitian setting. The quadratic $\epsilon$-pseudospectrum seems to be more helpful with finding localized states as described, with some added benefits. The construction of the quadratic composite operator is inspired by Cerjan, Loring and Vides in [cerjan_loring_vides_2023], where it is defined for systems described by Hermitian Hamiltonians. The quadratic composite operator first appears implicitly in [loring]. Because of the assumption taken in these prior studies that the constituent operators in the construction are Hermitian and the ‘probe site’, i.e., the choice of $(\bm{x},E)$ is real, it is guaranteed that the corresponding $\epsilon$-pseudospectrum is real i.e. $\bm{\lambda}$ in $\mathbb{R}^{n+1}$.
One key idea is that for $A$ Hermitian, $A\bm{\Psi}=0$ if and only if $A^{2}\bm{\Psi}=0$. Thus, it is reasonable to consider a similar procedure with the Clifford composite operator in [cerjan_koekebier_Schulz_Baldes]. In particular there is reason to believe that $A^{\dagger}A\bm{\Psi}=0$ can provide useful information about $A\bm{\Psi}=0$. We know that the singular values of the $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$ are the square roots of the eigenvalues of
@@
In particular this is useful since a small deviation from normality of the Clifford composite operator implies
@@
and small deviation from being Hermitian of the Clifford composite operator implies
@@
Thus we have that we can approximate the Clifford linear and radial gaps by the smallest singular value of the Clifford composite operator. By Equation 9 we have a way of defining a quadratic composite operator for the Clifford composite operator similar to the Hermitian setting in [cerjan_loring_vides_2023]. We do this by ignoring $F$ and the commutator terms in the computation of Equation 9. Specifically we could use
$\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I\right)+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}.$
Since the spectrum is determined by the diagonal block entries we could take one of them. Namely,
$\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}+(B-\nu)^{\dagger}(B-\nu).$
Note we don’t consider
$\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}+(B-\nu)(B-\nu)^{\dagger}.$
as the spectrum of both are the same with the exception of a zero eigenvalue. Instead of taking this as the definition of a quadratic composite operator, we will generalize this construction. We will allow various non-normal operators in the construction besides $B-\nu$. The proof that is to come about approximately localized eigenstates will follow with this more general construction.
@@
$\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=\left(\sum_{i=1}^{d_{1}}(A_{i}-\lambda_{i})^{2}\right)+\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}(B_{j}-\nu_{j})\right).$
Using this definition we have generalized the quadratic composite operator in [cerjan_loring_vides_2023]. In fact the following proposition generalizes Proposition II.1. in [cerjan_loring_vides_2023].
If we recall that a unit state $\bm{\psi}$ in a Hilbert space determines a probability distribution with respect to a normal matrix $M$. In particular for Hermitian matrices $M$ it is well established that we need only a complete basis for the eigenspace of $M$ to define a probablity distribution and as a result to define expectation and variance, see Appendix A of [hirvensalo_2001]. While we do not need to use the probability distribution directly, we do need the corresponding expectation and square variance for both Hermitian and non-Hermitian matrices. This allows us to talk loosely about approximate localization of a state in position, energy, momentum, frequency, etc. The mathematical notation that we will use for expectation and variance are as follows,
@@
    $\min_{||\bm{\psi}||=1}\sqrt{\sum_{i=1}^{d_{1}}\textup{V}^{2}_{\bm{\psi}}[A_{i}]+\left(\textup{E}_{\bm{\psi}}[A_{i}]-\lambda_{i}\right)^{2}+\sum_{j=1}^{d_{2}}\textup{V}^{2}_{\bm{\psi}}[B_{j}]+\left|\textup{E}_{\bm{\psi}}[B_{j}]-\nu_{j}\right|^{2}},$
@@
    $M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=\begin{bmatrix}A_{1}-\lambda_{1}\\ \vdots\\ A_{d_{1}}-\lambda_{d_{1}}\\ B_{1}-\nu_{1}\\ \vdots\\ B_{d_{2}}-\nu_{d_{2}}\end{bmatrix}.$
@@
    The square root for the smallest eigenvalue of $\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$.
@@
To show (1) $\Leftrightarrow$ (2) we have that $A$ is Hermitian and $\bm{\lambda}\in\mathbb{R}^{d_{1}}$ we have that $(A-\lambda_{i})^{2}=(A-\lambda_{i})^{\dagger}(A-\lambda_{i})$ and thus it is enough to prove the statement for $Q_{(\vec{\bm{0}},\bm{\nu})}(\bm{0},B)$.
$||B_{j}\bm{\psi}-\nu_{j}\bm{\psi}||^{2}$ $=\langle B_{j}\bm{\psi}-\nu_{j},B_{j}\bm{\psi}-\nu_{j}\rangle$
@@
$\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)^{\dagger}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right).$
To show (1) $\Leftrightarrow$ (3) we note by [golub_vanloan]
$\sigma_{\min}\left(M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)$ $=\min_{||\bm{\psi}||=1}\left\lVert M_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\bm{\psi}\right\rVert$
$=\min_{||\bm{\psi}||=1}\left\lVert\begin{bmatrix}A_{1}\bm{\psi}-\lambda_{1}\bm{\psi}\\ \vdots\\ A_{d_{1}}\bm{\psi}-\lambda_{d_{1}}\bm{\psi}\\ B_{1}\bm{\psi}-\nu_{1}\bm{\psi}\\ \vdots\\ B_{d_{2}}\bm{\psi}-\nu_{d_{2}}\bm{\psi}\end{bmatrix}\right\rVert$
$=\min_{||\bm{\psi}||=1}\sqrt{\sum_{i=1}^{d_{1}}||A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}||^{2}+\sum_{j=1}^{d_{2}}||B_{j}\bm{\psi}-\nu_{j}\bm{\psi}||^{2}}.$
Thus if we want a state that is approximately localized in position and energy we have reason to study the smallest eigenvalue of the quadratic composite operator.
@@
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian and probe site $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then the quadratic gap is defined as
$\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)=\sqrt{\sigma_{\min}\left(\text{Q}_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)}.$

### v2

@@ 3 Non-Hermitian spectral localizer
If the non-Hermitian spectral localizer is singular then we are done and $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. So we will assume that the non-Hermitian spectral localizer is invertible. We will use linearity to reduce to the case where $(\bm{\lambda},\nu)=\bm{0}$. If $2\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert B^{\dagger}B\bm{\psi}\right\rVert+\left\lVert BB^{\dagger}\bm{\psi}\right\rVert\leq\epsilon_{1}$ for some unit $\bm{\psi}$ we will define
$\bm{\Psi}=\frac{1}{\sqrt{2m}}\phi^{\oplus 2m}=\frac{1}{\sqrt{2m}}\begin{bmatrix}\bm{\psi}\\ \vdots\\ \bm{\psi}\end{bmatrix}$
@@
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert^{2}$ $=\frac{1}{2}\left[\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+B^{\dagger}B\bm{\psi}\right\rVert^{2}+\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+BB^{\dagger}\bm{\psi}\right\rVert^{2}\right]$
$\leq\frac{1}{2}\left[\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+B^{\dagger}B\bm{\psi}\right\rVert+\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+BB^{\dagger}\bm{\psi}\right\rVert\right]^{2}$
$\leq\frac{1}{2}\left[\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert B^{\dagger}B\bm{\psi}\right\rVert+\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert BB^{\dagger}\bm{\psi}\right\rVert\right]^{2}$
$\leq\frac{1}{2}\epsilon_{1}^{2}$
combined with Equation 11 that
$\left\lVert L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)^{\dagger}L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)\bm{\Psi}\right\rVert\leq\frac{1}{\sqrt{2}}\epsilon_{1}+\left(\sum_{i\not=k}\left\lVert[A_{i},A_{k}]\right\rVert\right)+\lVert F\rVert.$
Lastly Equation (12) finalizes the proof. ∎
Corollary 3.9.
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, $B$ in $\mathcal{M}_{n}(\mathbb{C})$ non-Hermitian. If there is a unit state $\bm{\psi}$ such that
$\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)$ $+\left\lVert B\bm{\psi}-\nu\bm{\psi}\right\rVert^{2}\leq\epsilon_{1}$
and
for all $i$, with $F$ as in Equation 10; then $(\bm{\lambda},\nu)$ is in $\dot{\Lambda}_{\epsilon}(\bm{A},B)$ with $\epsilon=\sqrt{\epsilon_{1}+\epsilon_{2}}$ or $(\bm{\lambda},\nu)$ is in $\dot{\Lambda}_{\epsilon=0}(\bm{A},B)$.
Proof.
If the non-Hermitian spectral localizer is singular then we are done and $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. So we will assume that the non-Hermitian spectral localizer is invertible. We will use linearity to reduce to the case where $(\bm{\lambda},\nu)=\bm{0}$. If $\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert B^{\dagger}B\bm{\psi}\right\rVert\leq\epsilon_{1}$ for some unit $\bm{\psi}$ we will define repeat the proof of Lemma (3.8), but instead defining
$\bm{\Psi}=\frac{1}{\sqrt{m}}\phi^{\oplus m}\oplus\bm{0}^{m}=\frac{1}{\sqrt{m}}\begin{bmatrix}\bm{\psi}\\ \vdots\\ \bm{\psi}\\ \bm{0}\\ \vdots\\ \bm{0}\end{bmatrix}$.
with $m$ as in Equation 8. Then by Equation 9, and $F$ as in Equation 10, we have
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+B^{\dagger}B\bm{\psi}\right\rVert^{2}$
$\leq\left[\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert B^{\dagger}B\bm{\psi}\right\rVert\right]^{2}$
$\leq\epsilon_{1}^{2}.$
The rest of the proof uses the techniques of the proof in Lemma (3.8). ∎
Corollary 3.10.
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian, $B$ in $\mathcal{M}_{n}(\mathbb{C})$ non-Hermitian. If there is a unit state $\bm{\psi}$ such that
$\left(\sum_{i=1}^{d}\left\lVert A_{i}\bm{\psi}-\lambda_{i}\bm{\psi}\right\rVert^{2}\right)$ $+\left\lVert B^{\dagger}\bm{\psi}-\nu^{*}\bm{\psi}\right\rVert^{2}\leq\epsilon_{1}$
and
for all $i$, with $F$ as in Equation 10; then $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon}(\bm{A},B)$ with $\epsilon=\sqrt{\epsilon_{1}+\epsilon_{2}}$ or $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. Here $\Lambda_{\epsilon}(\bm{A},B)$ can be either the radial or linear $\epsilon$-pseudospectrum.
Proof.
If the non-Hermitian spectral localizer is singular then we are done and $(\bm{\lambda},\nu)$ is in $\Lambda_{\epsilon=0}(\bm{A},B)$. So we will assume that the non-Hermitian spectral localizer is invertible. We will use linearity to reduce to the case where $(\bm{\lambda},\nu)=\bm{0}$. If $\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert BB^{\dagger}\bm{\psi}\right\rVert\leq\epsilon_{1}$ for some unit $\bm{\psi}$ we will define repeat the proof of Lemma (3.8), but instead defining
$\bm{\Psi}=\frac{1}{\sqrt{m}}\bm{0}^{m}\oplus\phi^{\oplus m}=\frac{1}{\sqrt{m}}\begin{bmatrix}\bm{0}\\ \vdots\\ \bm{0}\\ \bm{\psi}\\ \vdots\\ \bm{\psi}\end{bmatrix}$.
with $m$ as in Equation 8. Then by Equation 9, and $F$ as in Equation 10, we have
$\left\lVert\left(\left(\sum_{i=1}^{d}A_{i}^{2}\otimes I_{2m}\right)+C\right)\bm{\Psi}\right\rVert^{2}$ $=\left\lVert\sum_{i=1}^{d}A_{i}^{2}\bm{\psi}+BB^{\dagger}\bm{\psi}\right\rVert^{2}$
$\leq\left[\sum_{i=1}^{d}\left\lVert A_{i}^{2}\bm{\psi}\right\rVert+\left\lVert BB^{\dagger}\bm{\psi}\right\rVert\right]^{2}$
$\leq\epsilon_{1}^{2}$
The rest of the proof uses the techniques of the proof in Lemma (3.8). ∎
@@
The lack of a proof does not hinder us from proceeding with the generalization of the theory in [4]. The fact that for a normal matrix $A$,
@@
In any case we are working with a non-normal matrix, and with luck the matrix can be nearly Hermitian or nearly normal. However we do not work with luck and thus we need a way to characterize what we mean by nearly normal. We will make use of definition and terminology presented in [30] in order to characterize deviation from normality. In particular we have,
@@
This is in some sense equivalent to considering the quantity $AA^{\dagger}-A^{\dagger}A||$ and formal statements comparing the two quantities can be found in Theorem 2 of [10]. Also note that when the non-Hermitian spectral localizer does not deviate too much from normal then we get some decent spectral variation properties.
@@ 4 Quadratic composite operator
In [4] and [32], it was shown that when it comes to finding unit states that are approximate joint eigenvectors to the observables/operators, both the spectral localizer’s $\epsilon$-pseudospectrum and the quadratic’s $\epsilon$-pseudospectrum provide good methods in the setting when all observables are Hermitian. In Section 3 we have a non-Hermitian matrix/Hamiltonian in the construction of the non-Hermitian spectral localizer in order to generalize the spectral localizer $\epsilon$-pseudospectrum theory. The appeal of the spectral localizer $\epsilon$-pseudospectrum in the Hermitian case and non-Hermitian spectral localizer $\epsilon$-pseudospectrum in the non-Hermitian case is that it also provides information about a system’s topology see [6],[8],[9] for the Hermitian setting and [3] for the non-Hermitian setting. The quadratic $\epsilon$-pseudospectrum seems to be more helpful with finding localized states as described, with some added benefits. The construction of the quadratic composite operator is inspired by Cerjan, Loring and Vides in [4], where it is defined for systems described by Hermitian Hamiltonians. The quadratic composite operator first appears implicitly in [32]. Because of the assumption taken in these prior studies that the constituent operators in the construction are Hermitian and the ‘probe site’, i.e., the choice of $(\bm{x},E)$ is real, it is guaranteed that the corresponding $\epsilon$-pseudospectrum is real i.e. $\bm{\lambda}$ in $\mathbb{R}^{n+1}$.
One key idea is that for $A$ Hermitian, $A\bm{\Psi}=0$ if and only if $A^{2}\bm{\Psi}=0$. Thus, it is reasonable to consider a similar procedure with the non-Hermitian spectral localizer in [3]. It’s explicit and quick to prove that $A^{\dagger}A\bm{\Psi}=0$ if and only if $A\bm{\Psi}=0$. We know that the singular values of the $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$ are the square roots of the eigenvalues of
@@
In particular this is useful since a small deviation from normality of the non-Hermitian spectral localizer implies
@@
and small deviation from being Hermitian of the non-Hermitian spectral localizer implies
@@
Thus we have that we can approximate the Clifford linear and radial gaps by the smallest singular value of the non-Hermitian spectral localizer. By Equation 9 we have a way of defining a quadratic composite operator for the non-Hermitian spectral localizer similar to the Hermitian setting in [4]. We do this by ignoring $F$ and the commutator terms in the computation of Equation 9. Specifically we could use
$\left(\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I_{2m}\right)+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}I_{m}&\\ &0_{m}\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0_{m}&\\ &I_{m}\end{bmatrix}$.
Since the spectrum is determined by the diagonal block entries we could take the two distinct diagonal blocks. Namely,
$\sum_{i=1}^{d}(A_{i}-\lambda_{i})^{2}\otimes I_{2}+(B-\nu)^{\dagger}(B-\nu)\otimes\begin{bmatrix}1&\\ &0\end{bmatrix}+(B-\nu)(B-\nu)^{\dagger}\otimes\begin{bmatrix}0&\\ &1\end{bmatrix}$.
Instead of taking this as the definition of a quadratic composite operator, we will generalize this construction. We will allow various non-normal operators in the construction besides $B-\nu$. The proof that is to come about approximately localized eigenvectors will follow with this more general construction.
@@
$Q_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ $=RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\otimes\begin{bmatrix}1&\\ &0\end{bmatrix}+LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\otimes\begin{bmatrix}0&\\ &1\end{bmatrix}$.
With the right and left quadratic composite operators defined as
$RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ $=\left(\sum_{i=1}^{d_{1}}(A_{i}-\lambda_{i})^{2}\right)+\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})^{\dagger}(B_{j}-\nu_{j})\right)$
$LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ $=\left(\sum_{i=1}^{d_{1}}(A_{i}-\lambda_{i})^{2}\right)+\left(\sum_{j=1}^{d_{2}}(B_{j}-\nu_{j})(B_{j}-\nu_{j})^{\dagger}\right)$
Immediately from the definition we get the following Corollaries.
Corollary 4.2.
Suppose that we have tuples of matrices $\bm{A}\in\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),\bm{B}\in\Pi_{i=1}^{d_{2}}M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian, and $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then
$\sigma_{\min}\left(Q_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=\min\left\{\sigma_{\min}\left(RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right),\sigma_{\min}\left(LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)\right\}$
Note that when all the $B_{i}$ are normal, the spectrum of $LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ is equal to the spectrum of $RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ which gives us the following corollary.
Corollary 4.3.
Suppose that we have tuples of matrices $\bm{A}\in\Pi_{i=1}^{d_{1}}M_{n}(\mathbb{C}),\bm{B}\in\Pi_{i=1}^{d_{2}}M_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ normal. If $(\bm{\lambda},\bm{\nu})$ is in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then
$\sigma_{\min}\left(Q_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=\sigma_{\min}\left(RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=\sigma_{\min}\left(LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)$
Reflecting on the main point of Lemma 3.7 we expect that the eigenvector corresponding with the smallest eigenvalue of the quadratic operator will pick up either an approximate right or left eigenvector of $A_{i}$ and $B_{j}$. Using the definition of the quadratic composite operator above we have generalized the quadratic composite operator in [4]. In fact the following proposition generalizes Proposition II.1. in [4].
Recall that a unit state $\bm{\psi}$ in a Hilbert space determines a probability distribution with respect to a normal matrix $M$. In particular for Hermitian matrices $M$ it is well established that we need only a complete basis for the eigenspace of $M$ to define a probablity distribution and as a result to define expectation and variance, see Appendix A of [15]. While we do not need to use the probability distribution directly, we do need the corresponding expectation and square variance for both Hermitian and non-Hermitian matrices. This allows us to talk loosely about approximate localization of a state in position, energy, momentum, frequency, etc. The mathematical notation that we will use for expectation and variance are as follows,
@@
[…]
@@
    $RM_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=\begin{bmatrix}A_{1}-\lambda_{1}\\ \vdots\\ A_{d_{1}}-\lambda_{d_{1}}\\ B_{1}-\nu_{1}\\ \vdots\\ B_{d_{2}}-\nu_{d_{2}}\end{bmatrix}$.
@@
    The square root for the smallest eigenvalue of $RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$.
@@
[…]
@@
$\left(RM_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)^{\dagger}\left(RM_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)=RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right).$
To show (1) $\Leftrightarrow$ (3) we note by [12]
[…]
Remark 4.5.
A similar statement can be proven about the square root of the smallest eigenvalue of $LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)$ by noting that $LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)=RQ_{(\bm{\lambda},\bm{\nu}^{*})}\left(\bm{A},\bm{B}^{\dagger}\right)$.
Thus if we want a state that is approximately localized in position and energy we have reason to study the smallest eigenvalue of the right quadratic composite operator. If we want an approximate left eigenvector then we look at the smallest eigenvalue of the left quadratic composite operator.
@@
Suppose that $\bm{A}$ in $\Pi_{i=1}^{d_{1}}\mathcal{M}_{n}(\mathbb{C})$, $\bm{B}$ in $\Pi_{i=1}^{d_{2}}\mathcal{M}_{n}(\mathbb{C})$ with $A_{i}$ Hermitian, $B_{j}$ non-Hermitian and probe site $(\bm{\lambda},\bm{\nu})$ in $\mathbb{R}^{d_{1}}\oplus\mathbb{C}^{d_{2}}$, then the right quadratic gap is defined as
$\mu_{(\bm{\lambda},\bm{\nu})}^{\text{RQ}}\left(\bm{A},\bm{B}\right)=\sqrt{\sigma_{\min}\left(RQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)}.$
We can similarly define the left quadratic gap as being the square root of the smallest eigenvalue of the left quadratic composite operator and the quadratic gap function can be define as the square root of the smallest eigenvalue quadratic composite operator. They can be denoted as
$\mu_{(\bm{\lambda},\bm{\nu})}^{\text{LQ}}\left(\bm{A},\bm{B}\right)=\sqrt{\sigma_{\min}\left(LQ_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)}$
and
$\mu_{(\bm{\lambda},\bm{\nu})}^{\text{Q}}\left(\bm{A},\bm{B}\right)=\sqrt{\sigma_{\min}\left(Q_{(\bm{\lambda},\bm{\nu})}\left(\bm{A},\bm{B}\right)\right)}.$